### v1

@@ 1 Introduction
Large language models (LLMs) Brown et al. (2020); Touvron et al. (2023); Achiam et al. (2024) are renowned for encapsulating a vast repository of world knowledge Petroni et al. (2019); Roberts et al. (2020), referred to as parametric knowledge. These models excel in knowledge-intensive tasks including QA Petroni et al. (2019), fact-checking Gao et al. (2023a), knowledge generation Chen et al. (2023c), inter alia. In the meantime, LLMs continue to engage with external contextual knowledge after deployed Pan et al. (2022), including user prompts Liu et al. (2023a), interactive dialogues Zhang et al. (2020), or retrieved documents from the Web Lewis et al. (2020); Shi et al. (2023c), and tools Schick et al. (2023); Zhuang et al. (2023).
@@
Knowledge conflict is originally rooted in open-domain QA research. The concept gained attention in Longpre et al. (2021) that focused on the entity-based conflicts between parametric knowledge and external passages. Concurrently, discrepancies among multiple passages were also scrutinized subsequently Chen et al. (2022). Knowledge conflicts attract significant attention with the recent advent of LLMs. For instance, recent studies find that LLMs exhibit both adherence to parametric knowledge and susceptibility to contextual influences Xie et al. (2023), which can be problematic when this external knowledge is factually incorrect Pan et al. (2023b). Given the implications for the trustworthiness (Du et al., 2022b), real-time accuracy (Kasai et al., 2022), and robustness of LLMs (Ying et al., 2023), it is imperative to delve deeper into understanding and resolving knowledge conflicts (Xie et al., 2023; Wang et al., 2023g).
@@
The methodology of our survey is illustrated in Figure 2, we conceptualize lifecycle of knowledge conflicts as both a cause leading to various behaviors in the model, and an effect emerges from the intricate nature of knowledge. Knowledge conflicts serve as a crucial intermediary between causes and model behaviors. For instance, they significantly contribute to the model generating factually incorrect information, a.k.a., hallucinations Ji et al. (2023); Zhang et al. (2023d). Our research, in a manner akin to Freudian psychoanalysis, underscores the significance of understanding the origins of these conflicts. Although existing analyses Chen et al. (2022); Xie et al. (2023); Wang et al. (2023g) tend to construct such conflicts artificially, we posit that these analyses do not sufficiently address the interconnectedness of the issue.
@@
(§ 2.2) , color=lightcoral!100, fill=lightcoral!15, thick, text=black [ Open-domain QA , color=lightcoral!100, fill=lightcoral!15, thick, text=black [ Longpre et al. (2021), Chen et al. (2022), Tan et al. (2024) , leaf, text width = 12em, color=lightcoral!100, fill=lightcoral!15, thick, text=black ] ] [ General , color=lightcoral!100, fill=lightcoral!15, thick, text=black [ Xie et al. (2023), Wang et al. (2023g), Ying et al. (2023),
@@
of Knowledge , color=cyan!100, fill=cyan!15, thick, text=black [ Chuang et al. (2023), Li et al. (2023c) , leaf,color=cyan!100, fill=cyan!15, thick, text=black, text width=12em ] ] [ Cross-lingual Inconsistency , color=cyan!100, fill=cyan!15, thick, text=black [ Wang et al. (2023e), Qi et al. (2023) ,leaf,color=cyan!100, fill=cyan!15, thick, text=black, text width=12em ] ] ] [ Solution
@@
Figure 3: Taxonomy of knowledge conflicts. We mainly list works in the era of large language models. Refer to caption denotes pre-hoc solution and Refer to caption denotes post-hoc solution.
Going beyond reviewing and analyzing causes and behaviors, we delve deeper to provide a systematic review of solutions, which are employed to minimize the undesirable consequences of knowledge conflicts, i.e., to encourage the model to exhibit desired behaviors that conform to specific objectives (please noted that these objectives may differ based on the particular scenario). Based on the timing relative to potential conflicts, strategies are divided into two categories: pre-hoc and post-hoc strategies. The key distinction between them lies in whether adjustments are made before or after potential conflicts arise***Another interpretation is that a pre-hoc strategy is proactive while a post-hoc one is reactive.. The taxonomy of knowledge conflicts is outlined in Figure 3. We sequentially discuss the three kinds of knowledge conflicts, detailing for each the causes, analysis of model behaviors, and available solutions organized according to their respective objectives. Related datasets can be found in Table 1.
@@
Xie et al. (2023)
Gen
PopQA (2023), StrategyQA (Geva et al. (2021))
20,091
CM3
KC (2023g)
Sub
N/A (LLM generated)
9,803
CM
KRE (2023)
Gen
MuSiQue (2022), SQuAD2.0 (2018), ECQA (2021), e-CARE (2022a)
11,684
CM
Farm (2023)
Gen
BoolQ (2019), NQ (2019), TruthfulQA (2022)
1,952
CM
Tan et al. (2024)
Gen
NQ (2019), TriviaQA (2017)
14,923
CM
WikiContradiction (2021)
Hum
Wikipedia
2,210
IC
ClaimDiff (2022)
Hum
N/A
2,941
IC
Pan et al. (2023a)
Gen,Sub
SQuAD v1.1 (2016)
52,189
IC
ContraDoc (2023a)
Gen
CNN-DailyMail (2015), NarrativeQA (2018), WikiText (2017)
449
IC
ConflictingQA (2024)
Gen
N/A
238
IC
ParaRel (2021)
Hum
T-REx (2018)
328
IM
@@
  • 3. [Uncaptioned image] For CM datasets, conflicts are derived from a certain model’s parametric knowledge, which can vary between models. Therefore, one should select a subset of the dataset that aligns with the tested model’s knowledge when using CM datasets.
Table 1: Datasets on evaluating a large language model’s behavior when encountering knowledge conflicts. CM: context-memory conflict, IC: inter-context conflict, IM: intra-memory conflict.
@@ 2.1 Causes
Temporal Misalignment. Temporal misalignment naturally arises in models trained on data collected in the past, as they may not accurately reflect contemporary or future realities (i.e., the contextual knowledge after the deployment) Luu et al. (2021); Lazaridou et al. (2021); Liska et al. (2022). Such misalignment can degrade the model’s performance and relevancy over time, as it may fail to capture new trends, shifts in language use, cultural changes, or updates in knowledge. Researchers have noted that temporal misalignment relegates the model’s performance on various NLP tasks Luu et al. (2021); Zhang and Choi (2021); Dhingra et al. (2022); Kasai et al. (2022); Cheang et al. (2023). Furthermore, the issue of temporal misalignment is expected to intensify due to the pre-training paradigm and the escalating costs associated with scaling up models Chowdhery et al. (2023); Achiam et al. (2024).
@@
Prior work tries to tackle temporal misalignment by focusing on three lines of strategies: Knowledge editing (KE) aims to directly update the parametric knowledge of an existing pre-trained model (Sinitsin et al., 2020; De Cao et al., 2021; Mitchell et al., 2021; Onoe et al., 2023). Retrieval-augmented generation (RAG) leverages a retrieval module to fetch relevant documents from external sources (e.g., database, the Web) to supplement the model’s knowledge without altering its parameters Karpukhin et al. (2020); Guu et al. (2020); Lewis et al. (2020); Lazaridou et al. (2022); Borgeaud et al. (2022); Peng et al. (2023); Vu et al. (2023). Continue learning (CL) seeks to update the internal knowledge through continual pre-training on new and updated data Lazaridou et al. (2021); Jang et al. (2021, 2022). However, these methods of mitigating temporal misalignment are not magic bullets. KE can bring in side effects of knowledge conflict, leading to knowledge inconsistency (i.e., a sort of intra-memory conflict) and may even enhance the hallucination of LLMs Li et al. (2023f); Pinter and Elhadad (2023). For RAG, it is inevitable to encounter knowledge conflicts since the model’s parameters are not updated Chen et al. (2021); Zhang and Choi (2021). CL suffers from catastrophic forgetting issues and demands significant computational resources De Lange et al. (2021); He et al. (2021); Wang et al. (2023f).
@@
Misinformation Pollution. Misinformation pollution emerges as another contributor to context-memory conflict, particularly for time-invariant knowledge Jang et al. (2021) that the model has accurately learned. Adversaries exploit this vulnerability by introducing false or misleading information into the Web corpus of retrieved documents Pan et al. (2023a, b); Weller et al. (2022) and user conversations Xu et al. (2023). The latter poses a practical threat, as adversaries can leverage techniques such as prompt injection attacks Liu et al. (2023b); Greshake et al. (2023); Yi et al. (2023). This vulnerability poses a real threat, as models might unknowingly propagate misinformation if they incorporate deceptive inputs without scrutiny Xie et al. (2023); Pan et al. (2023b); Xu et al. (2023).
@@
Researchers observe that fabricated, malicious misinformation can markedly undermine the accuracy of automated fact-checking systems Du et al. (2022b) and open-domain question-answering models Pan et al. (2023a, b). Furthermore, recent studies also highlight the model’s tendency to align with user opinions, a.k.a., sycophancy, further exacerbating the issue Perez et al. (2022); Turpin et al. (2023); Wei et al. (2023); Sharma et al. (2023). In the current landscape of LLMs, there is growing apprehension in the NLP community regarding the potential generation of misinformation by LLMs Ayoobi et al. (2023); Kidd and Birhane (2023); Carlini et al. (2023); Zhou et al. (2023c); Spitale et al. (2023); Chen and Shu (2023b). Researchers acknowledge the challenges associated with detecting misinformation generated by LLMs Tang et al. (2023); Chen and Shu (2023a); Jiang et al. (2023). This underscores the urgency of addressing the nuanced challenges posed by LLMs in the context of contextual misinformation.
@@ 2.2 Analysis of Model Behaviors
General. Xie et al. (2023) leverage LLMs to generate conflicting context alongside the memorized knowledge. They find that LLMs are highly receptive to external evidence, even when it conflicts with their parametric, provided that the external knowledge is coherent and convincing. Meanwhile, they also identify a strong confirmation bias Nickerson (1998) in LLMs, i.e., the models tend to favor information consistent with their internal memory, even when confronted with conflicting external evidence. Wang et al. (2023g) posit that the desired behaviors when an LLM encounters conflicts should be to pinpoint the conflicts and provide distinct answers. They find while LLMs perform well in identifying the existence of knowledge conflicts, they struggle to determine the specific conflicting segments and produce a response with distinct answers amidst conflicting information. Ying et al. (2023) analyze the robustness of LLMs under conflicts with a focus on two perspectives: factual robustness (the ability to identify correct facts from prompts or memory) and decision style (categorizing LLMs’ behavior as intuitive, dependent, or rational-based on cognitive theory). The study finds that LLMs are highly susceptible to misleading prompts, especially in the context of commonsense knowledge. Qian et al. (2023) evaluate the potential interaction between parametric and external knowledge more systematically, cooperating knowledge graph (KG). They reveal that LLMs often deviate from their parametric knowledge when presented with direct conflicts or detailed contextual changes. Xu et al. (2023) study how large language models (LLMs) respond to knowledge conflicts during interactive sessions. Their findings suggest LLMs tend to favor logically structured knowledge, even when it contradicts factual accuracy.
@@ 2.3 Solutions
Faithful to Context. Fine-tuning. Li et al. (2022a) argue that an LLM should prioritize context for task-relevant information and rely on internal knowledge when the context is unrelated. They name the two properties controllability and robustness. They introduce Knowledge Aware FineTuning (KAFT) to strengthen the two properties by incorporating counterfactual and irrelevant contexts to standard training datasets. Gekhman et al. (2023) introduce TrueTeacher, which focuses on improving factual consistency in summarization by annotating model-generated summaries with LLMs. This approach helps in maintaining faithfulness to the context of the original documents, ensuring that generated summaries remain accurate without being misled by irrelevant or incorrect details. DIAL (Xue et al., 2023) focuses on improving factual consistency in dialogue systems via direct knowledge enhancement and RLFC for aligning responses accurately with provided factual knowledge.
@@
Decoding. Shi et al. (2023a) introduce Context-aware Decoding (CAD) to reduce hallucinations by amplifying the difference in output probabilities with and without context, which is similar to the concept of contrastive decoding Li et al. (2022c). CAD enhances faithfulness in LLMs by effectively prioritizing relevant context over the model’s prior knowledge, especially in tasks with conflicting information.
Knowledge Plug-in. Lee et al. (2022a) propose Continuously-updated QA (CuQA) for improving LMs’ ability to integrate new knowledge. Their approach uses plug-and-play modules to store updated knowledge, ensuring the original model remains unaffected. Unlike traditional continue pre-training or fine-tuning approaches, CuQA can solve knowledge conflicts.
@@
Query Augmentation. Weller et al. (2022) leverage the redundancy of information in large corpora to defend misinformation pollution. Their method involves query augmentation to find a diverse set of less likely poisoned passages, coupled with a confidence method named Confidence from Answer Redundancy (CAR), which compares the predicted answer’s consistency across retrieved contexts. This strategy mitigates knowledge conflicts by ensuring the model’s faithfulness through the cross-verification of answers in multiple sources.
@@
Disentangling Sources. DisentQA (Neeman et al., 2022) trains a model that predicts two types of answers for a given question: one based on contextual knowledge and one on parametric knowledge. Wang et al. (2023g) introduce a method to improve LLMs’ handling of knowledge conflicts. Their approach is a three-step process designed to help LLMs detect conflicts, accurately identify the conflicting segments, and generate distinct, informed responses based on the conflicting data, aiming for more precise and nuanced model outputs.
[…]
Remarks. Current mitigation approaches have contradicted goals because they do not distinguish between the two causes of knowledge conflict when considering conflict scenarios. Blindly being “faithful” to context or knowledge is undesirable. Some researchers regard that LLM should not rely solely on either parametric or contextual information but instead grant LLM users the agency to make informed decisions based on distinct answers (Wang et al., 2023g; Floridi, 2023).

### v2

@@ 1 Introduction
Large language models (LLMs; Brown et al. 2020; Touvron et al. 2023; OpenAI 2024) are renowned for encapsulating a vast repository of world knowledge Roberts et al. (2020); Hu et al. (2023), referred to as parametric knowledge. These models excel in knowledge-intensive tasks including QA Petroni et al. (2019), fact-checking Gao et al. (2023a), dialogue system Wang et al. (2023e), knowledge generation Chen et al. (2023c), inter alia. In the meantime, LLMs continue to engage with external contextual knowledge after deployed Pan et al. (2022), including user prompts Liu et al. (2023a), interactive dialogues Zhang et al. (2020); Wang et al. (2024a), or retrieved documents from the Web Lewis et al. (2020); Shi et al. (2023c), and tools Schick et al. (2023); Zhuang et al. (2023).
@@
Knowledge conflict is originally rooted in open-domain QA research. The concept gained attention in Longpre et al. (2021) that focused on the entity-based conflicts between parametric knowledge and external passages. Concurrently, discrepancies among multiple passages were also scrutinized subsequently Chen et al. (2022). Knowledge conflicts attract significant attention with the recent advent of LLMs. For instance, recent studies find that LLMs exhibit both adherence to parametric knowledge and susceptibility to contextual influences Xie et al. (2023), which can be problematic when this external knowledge is factually incorrect Pan et al. (2023b). Given the implications for the trustworthiness (Du et al., 2022b), real-time accuracy (Kasai et al., 2022), and robustness of LLMs (Ying et al., 2023), it is imperative to delve deeper into understanding and resolving knowledge conflicts (Xie et al., 2023; Wang et al., 2023h).
@@
We conceptualize the lifecycle of knowledge conflicts as both a cause leading to various behaviors, and an effect emerges from the intricate nature of knowledge as in Figure 2. Knowledge conflicts serve as a crucial intermediary between causes and model behaviors. For instance, they significantly contribute to the model generating factually incorrect information, a.k.a., hallucinations Ji et al. (2023); Zhang et al. (2023d). Our research, in a manner akin to Freudian psychoanalysis, underscores the significance of understanding the origins of these conflicts. Although existing analyses Chen et al. (2022); Xie et al. (2023); Wang et al. (2023h) tend to construct such conflicts artificially, we posit that these analyses do not sufficiently address the interconnectedness of the issue.
@@
(§ 2.2) , color=lightcoral!100, fill=lightcoral!15, thick, text=black [ Open-domain QA , color=lightcoral!100, fill=lightcoral!15, thick, text=black [ Longpre et al. (2021), Chen et al. (2022), Tan et al. (2024) , leaf, text width = 12em, color=lightcoral!100, fill=lightcoral!15, thick, text=black ] ] [ General , color=lightcoral!100, fill=lightcoral!15, thick, text=black [ Xie et al. (2023), Wang et al. (2023h), Ying et al. (2023),
@@
of Knowledge , color=cyan!100, fill=cyan!15, thick, text=black [ Chuang et al. (2023), Li et al. (2023c) , leaf,color=cyan!100, fill=cyan!15, thick, text=black, text width=12em ] ] [ Cross-lingual Inconsistency , color=cyan!100, fill=cyan!15, thick, text=black [ Wang et al. (2023f), Qi et al. (2023) ,leaf,color=cyan!100, fill=cyan!15, thick, text=black, text width=12em ] ] ] [ Solution
@@
Figure 3: Taxonomy of knowledge conflicts. We mainly list works in the era of LLMs. Refer to caption denotes pre-hoc solution and Refer to caption denotes post-hoc solution.
Going beyond reviewing and analyzing causes and behaviors, we delve deeper to provide a systematic review of solutions, which are employed to minimize the undesirable consequences of knowledge conflicts, i.e., to encourage the model to exhibit desired behaviors that conform to specific objectives (please noted that these objectives may differ based on the particular scenario). Based on the timing relative to potential conflicts, strategies are divided into two categories: pre-hoc and post-hoc strategies. The key distinction between them lies in whether adjustments are made before or after potential conflicts arise [^1] Another interpretation is that a pre-hoc strategy is proactive while a post-hoc one is reactive.. The taxonomy of knowledge conflicts is outlined in Figure 3. We sequentially discuss the three kinds of knowledge conflicts, detailing for each the causes, analysis of model behaviors, and available solutions organized according to their respective objectives. Related datasets can be found in Table 1.
@@
Xie et al. (2023) Gen PopQA (2023), StrategyQA (Geva et al. (2021)) 20,091 CM3
KC (2023h) Sub N/A (LLM generated) 9,803 CM
KRE (2023) Gen MuSiQue (2022), SQuAD2.0 (2018), ECQA (2021), e-CARE (2022a) 11,684 CM
Farm (2023) Gen BoolQ (2019), NQ (2019), TruthfulQA (2022) 1,952 CM
Tan et al. (2024) Gen NQ (2019), TriviaQA (2017) 14,923 CM
WikiContradiction (2021) Hum Wikipedia 2,210 IC
ClaimDiff (2022) Hum N/A 2,941 IC
Pan et al. (2023a) Gen,Sub SQuAD v1.1 (2016) 52,189 IC
ContraDoc (2023a) Gen CNN-DailyMail (2015), NarrativeQA (2018), WikiText (2017) 449 IC
ConflictingQA (2024) Gen N/A 238 IC
ParaRel (2021) Hum T-REx (2018) 328 IM
@@
  • 3. [Uncaptioned image] When using CM datasets, conflicts arise from the specific model’s parametric knowledge, which can differ across models. Therefore, selecting a subset of the dataset that aligns with the tested model’s knowledge is crucial.
Table 1: Datasets on evaluating LLMs’ behavior when encountering knowledge conflicts. CM: context-memory conflict, IC: inter-context conflict, IM: intra-memory conflict.
@@ 2.1 Causes
Temporal Misalignment. Temporal misalignment naturally arises in models trained on data collected in the past, as they may not accurately reflect contemporary or future realities (i.e., the contextual knowledge after the deployment) Luu et al. (2021); Lazaridou et al. (2021); Liska et al. (2022). Such misalignment can degrade the model’s performance and relevancy over time, as it may fail to capture new trends, shifts in language use, cultural changes, or updates in knowledge. Researchers have noted that temporal misalignment relegates the model’s performance on various NLP tasks Luu et al. (2021); Zhang and Choi (2021); Dhingra et al. (2022); Kasai et al. (2022); Cheang et al. (2023). Furthermore, the issue of temporal misalignment is expected to intensify due to the pre-training paradigm and the escalating costs associated with scaling up models (Kaplan et al., 2020).
@@
Prior work tries to tackle temporal misalignment by focusing on three lines of strategies: Knowledge editing (KE) aims to directly update the parametric knowledge of an existing pre-trained model (Sinitsin et al., 2020; De Cao et al., 2021; Mitchell et al., 2021; Onoe et al., 2023). Retrieval-augmented generation (RAG) leverages a retrieval module to fetch relevant documents from external sources (e.g., database, the Web) to supplement the model’s knowledge without altering its parameters Karpukhin et al. (2020); Guu et al. (2020); Lewis et al. (2020); Lazaridou et al. (2022); Borgeaud et al. (2022); Peng et al. (2023); Vu et al. (2023). Continue learning (CL) seeks to update the internal knowledge through continual pre-training on new and updated data Lazaridou et al. (2021); Jang et al. (2021, 2022). However, these methods of mitigating temporal misalignment are not magic bullets. KE can bring in side effects of knowledge conflict, leading to knowledge inconsistency (i.e., a sort of intra-memory conflict) and may even enhance the hallucination of LLMs Li et al. (2023f); Pinter and Elhadad (2023). For RAG, it is inevitable to encounter knowledge conflicts since the model’s parameters are not updated Chen et al. (2021); Zhang and Choi (2021). CL suffers from catastrophic forgetting issues and demands significant computational resources De Lange et al. (2021); He et al. (2021); Wang et al. (2023g).
@@
Misinformation Pollution. Misinformation pollution emerges as another contributor to context-memory conflict, particularly for time-invariant knowledge Jang et al. (2021). Adversaries exploit this vulnerability by introducing false or misleading information into the Web corpus of retrieved documents Pan et al. (2023a, b); Weller et al. (2022) and user conversations Xu et al. (2023); Hu et al. (2024). The latter poses a practical threat, as adversaries can leverage techniques such as prompt injection attacks Liu et al. (2023b); Greshake et al. (2023); Yi et al. (2023). This vulnerability poses a real threat, as models might unknowingly propagate misinformation if they incorporate deceptive inputs without scrutiny Xie et al. (2023); Pan et al. (2023b); Xu et al. (2023).
@@
Fabricated, malicious misinformation can markedly undermine the accuracy of automated fact-checking Du et al. (2022b) and open-domain question-answering systems Pan et al. (2023a, b). Furthermore, recent studies also highlight the model’s tendency to align with user opinions, a.k.a., sycophancy, further exacerbating the issue Perez et al. (2022); Turpin et al. (2023); Wei et al. (2023); Sharma et al. (2023). In the current landscape of LLMs, there is growing apprehension in the NLP community regarding the potential generation of misinformation by LLMs Ayoobi et al. (2023); Kidd and Birhane (2023); Carlini et al. (2023); Zhou et al. (2023c); Spitale et al. (2023); Chen and Shu (2023b). Researchers acknowledge the challenges associated with detecting misinformation generated by LLMs Tang et al. (2023); Chen and Shu (2023a); Jiang et al. (2023). This underscores the urgency of addressing the nuanced challenges posed by LLMs in the context of contextual misinformation.
@@ 2.2 Analysis of Model Behaviors
General. Xie et al. (2023) leverage LLMs to generate conflicting context alongside the memorized knowledge. They find that LLMs are highly receptive to external evidence, even when it conflicts with their parametric, provided that the external knowledge is coherent and convincing. Meanwhile, they also identify a strong confirmation bias Nickerson (1998) in LLMs, i.e., the models tend to favor information consistent with their internal memory, even when confronted with conflicting external evidence. Wang et al. (2023h) posit that the desired behaviors when an LLM encounters conflicts should be to pinpoint the conflicts and provide distinct answers. While LLMs perform well in identifying the existence of knowledge conflicts, they struggle to determine the specific conflicting segments and produce a response with distinct answers amidst conflicting information. Ying et al. (2023) analyze the robustness of LLMs under conflicts with a focus on two perspectives: factual robustness (the ability to identify correct facts from prompts or memory) and decision style (categorizing LLMs’ behavior as intuitive, dependent, or rational-based on cognitive theory). The study finds that LLMs are highly susceptible to misleading prompts, especially in the context of commonsense knowledge. Qian et al. (2023) evaluate the potential interaction between parametric and external knowledge more systematically, cooperating knowledge graph (KG). They reveal that LLMs often deviate from their parametric knowledge when presented with direct conflicts or detailed contextual changes. Xu et al. (2023) study how LLMs respond to knowledge conflicts during interactive sessions. Their findings suggest LLMs tend to favor logically structured knowledge, even when it contradicts factual accuracy.
@@ 2.3 Solutions
Faithful to Context. Fine-tuning. Li et al. (2022a) argue that an LLM should prioritize context for task-relevant information and rely on internal knowledge when the context is unrelated. They name the two properties controllability and robustness. They introduce Knowledge Aware FineTuning (KAFT) to strengthen the two properties by incorporating counterfactual and irrelevant contexts into standard training datasets. TrueTeacher (Gekhman et al., 2023) focuses on improving factual consistency in summarization by annotating model-generated summaries with LLMs. This approach helps in maintaining faithfulness to the context of the original documents, ensuring that generated summaries remain accurate without being misled by irrelevant or incorrect details. DIAL (Xue et al., 2023) improves factual consistency in dialogue systems via direct knowledge enhancement and reinforcement learning for factual consistency (RLFC) for aligning responses accurately with provided factual knowledge.
@@
Decoding. Shi et al. (2023a) introduce Context-aware Decoding (CAD) to reduce hallucinations by amplifying the difference in output probabilities with and without context, similar to the concept of contrastive decoding Li et al. (2022c). CAD enhances faithfulness in LLMs by prioritizing relevant context over the model’s prior knowledge, especially in tasks with conflicting information.
Knowledge Plug-in. Lee et al. (2022a) propose Continuously-updated QA (CuQA) for improving LMs’ ability to integrate new knowledge. Their approach uses plug-and-play modules to store updated knowledge, ensuring the original model remains unaffected. Unlike traditional continued pre-training or fine-tuning approaches, CuQA can solve knowledge conflicts.
@@
Query Augmentation. Weller et al. (2022) leverage the redundancy of information in large corpora to defend misinformation pollution. Their method involves query augmentation to find a diverse set of less likely poisoned passages, coupled with a confidence method named Confidence from Answer Redundancy, which compares the predicted answer’s consistency across retrieved contexts. This strategy mitigates knowledge conflicts by ensuring the model’s faithfulness through the cross-verification of answers from multiple sources.
@@
Disentangling Sources. DisentQA (Neeman et al., 2022) trains a model that predicts two types of answers for a given question: one based on contextual knowledge and one on parametric knowledge. Wang et al. (2023h) introduce a method to improve LLMs’ handling of knowledge conflicts. Their approach is a three-step process designed to help LLMs detect conflicts, accurately identify the conflicting segments, and generate distinct, informed responses based on the conflicting data, aiming for more precise and nuanced model outputs.
Improving Factuality. Zhang et al. (2023e) propose COMBO, a framework that pairs compatible generated and retrieved passages to resolve discrepancies. It uses discriminators trained on silver labels to assess passage compatibility, improving ODQA performance by leveraging both LLM-generated (parametric) and external retrieved knowledge. Jin et al. (2024a) introduce a contrastive-decoding-based algorithm, namely CD2, which maximizes the difference between various logits under knowledge conflicts and calibrates the model’s confidence in the truthful answer.
Remarks. Current mitigation approaches have contradicted goals because they do not distinguish between the two causes of knowledge conflict when considering conflict scenarios. Blindly being “faithful” to context or knowledge is undesirable. Some researchers regard that LLM should not rely solely on either parametric or contextual information but instead grant LLM users the agency to make informed decisions based on distinct answers (Wang et al., 2023h; Floridi, 2023).